### v1

@@ 4 Quasi-geodesics with respect to the generalized canonical metric
A geodesic on a manifold is a smooth curve with vanishing acceleration. In the dawn of Riemannian optimization, the solution is updated by searching along the geodesic which goes through the current solution in an appropriate tangent direction EdelAS98 ; Gaba82 ; Smit94 . In the Euclidean space, zero acceleration means $\ddot{X}(t)=0$ which implies that $X(t)=\dot{X}(0)t+X(0)$ which is a straight line. That is, geodesics can be viewed as a generalization of lines in the Euclidean space to the Riemannian manifold. In a general setting, the notion of acceleration depends on the Riemannian connection (AbsiMS08, , Sect. 5.4) and therefore the differential equation determining geodesics is usually much more involved. This makes the task of finding an explicit formula for the geodesic, given initial information $X(0),\dot{X}(0)$, very difficult. Fortunately, in optimization, only local information encoded by geodesics and its feasibility are interesting due to their usage in updating the iterate. Thus, one usually retains these features in approximating geodesics which results in the so-called quasi-geodesics, see, e.g., JurdML2020 ; KrakMLB2017 ; NishA2005 for more approaches to the quasi-geodesics on the Stiefel manifold.
In this paper, we follow a procedure similar to that of (GSAS21, , Subsect. 4.3), which was adopted from (EdelAS98, , Subsect. 2.2). Namely, we replace the acceleration that should be formulated in accordance with the Riemannian connection associated with the proposed generalized canonical metric by the one corresponding to the Euclidean metric, which turns out to be $\ddot{Y}(t).$ In view of the normal vectors formulated in Proposition 1, we define the quasi-geodesic with respect to the generalized canonical metric $g_{\rho,X_{\perp}}$ to be any solution $Y_{\rho}(t)\in\mathrm{iSt}(k,n)$ of the initial value problem
@@
with some $U(t)$ such that $JU\in\mathrm{Sym}(k)$ for all $t\geq 0$, although the obtained result apparently holds for $t<0$ too. In this section, a closed-form solution to this equation is presented. For the sake of brevity, we refrain from writing the variable $t$ in our presentation.
@@
$(-YU)^{T}AY+2\dot{Y}^{T}A\dot{Y}+Y^{T}A(-YU)=0$ (28)
or $\dot{Y}^{T}A\dot{Y}=JU$ due to the symmetry of $JU$. Hence, $U(t)=J\dot{Y}^{T}A\dot{Y}$ and therefore, the equation (25) becomes
@@
Next, let us set $W\coloneqq Y^{T}A\dot{Y}$ and $V\coloneqq\dot{Y}^{T}A\dot{Y}$. Thanks to the relation (26), $W\in\mathrm{Skew}(k)$ for $t\geq 0.$ Moreover, using (29), we get
$\dot{W}=V+Y^{T}A\ddot{Y}=V-Y^{T}AYJV=V-V=0,\mbox{ for }t\geq 0.$
Hence, $W(t)=W(0)\eqqcolon W_{0}\in\mathrm{Skew}(k)$ for $t\geq 0$. We proceed to determine $V$ which is symmetric. Taking the derivative of $V$ with (29) in mind, we have that
@@
$\dot{V}=W_{0}JV-VJW_{0},\ V(0)=V_{0},$ (30)
admits a solution $V(t)=\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}.$ Indeed, differentiating such $V(t)$ yields
$\dot{V}$ $=-\mathrm{e}^{-tW_{0}^{T}J}W_{0}^{T}JV_{0}\mathrm{e}^{-tJW_{0}}-\mathrm{e}^{-tW_{0}^{T}J}V_{0}JW_{0}\mathrm{e}^{-tJW_{0}}$
$=-W_{0}^{T}J\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}-\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}JW_{0}$
$=-W_{0}^{T}JV-VJW_{0}=W_{0}JV-VJW_{0}.$
Here, we invoke the skew-symmetry of $W_{0}$ and the identity $M\mathrm{e}^{tM}=\mathrm{e}^{tM}M$ for any real square matrix $M$ and any real number $t$.
Next, we present a technical lemma, which is similar to (GSAS21, , Lem. 4.6).
@@
$Y_{\rho}(t)=Y_{\rho}^{\mathrm{qgeo}}(t,Z)\coloneqq[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}},$ (33)
where $W_{0}=Y_{\rho}^{T}(0)A\dot{Y}_{\rho}(0)=X^{T}AZ$ and $V_{0}\coloneq\dot{Y}_{\rho}^{T}(0)A\dot{Y}_{\rho}(0)=Z^{T}AZ.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(Y_{\rho}(t)\mathrm{e}^{tJW_{0}}\right)$ $=\dot{Y_{\rho}}(t)\mathrm{e}^{tJW_{0}}+Y_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0}.$
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}\right)$ $=\ddot{Y_{\rho}}(t)\mathrm{e}^{tJW_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0}$
$=\left(-Y_{\rho}(t)JV\right)\mathrm{e}^{tJW_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0}$
$=-Y_{\rho}(t)J\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}\mathrm{e}^{tJW_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0}$
$=-Y_{\rho}(t)\mathrm{e}^{tJW_{0}}JV_{0}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}JW_{0},$
in which we have invoked (30) to replace $V\coloneq\dot{Y_{\rho}}^{T}A\dot{Y_{\rho}}$ by its explicit form $V=\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}$ when $Y_{\rho}(t)\in\mathrm{iSt}(k,n),$ for all $t\geq 0.$ Therefore,
$\frac{\mathrm{d}}{\mathrm{d}t}\left[Y_{\rho}(t)\mathrm{e}^{tJW_{0}}\;\;\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}\right]=\left[Y_{\rho}(t)\mathrm{e}^{tJW_{0}}\;\;\dot{Y}_{\rho}(t)\mathrm{e}^{tJW_{0}}\right]\begin{bmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{bmatrix}$. (34)
@@
$Y_{\rho}(t)$ $=\left[Y_{\rho}(0)\;\;\dot{Y}_{\rho}(0)\right]\mathrm{exp}\left(t\begin{bmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}},$
@@
Next, we will show that this is indeed a global solution to the problem (32). For brevity, we set $\Psi\coloneq\left[\begin{smallmatrix}JW_{0}&-JV_{0}\\ I_{k}&JW_{0}\end{smallmatrix}\right]$. In view of (33), it holds that
$\dot{Y}_{\rho}\mathrm{e}^{tJW_{0}}+Y_{\rho}\mathrm{e}^{tJW_{0}}JW_{0}=\frac{\mathrm{d}}{\mathrm{d}t}\left(Y_{\rho}\mathrm{e}^{tJW_{0}}\right)=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}I_{k}\\ 0\end{bmatrix}$.
@@
$\dot{Y}_{\rho}\mathrm{e}^{tJW_{0}}=$ $[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}I_{k}\\ 0\end{bmatrix}-[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}JW_{0}$ (35)
$=$ $[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\biggr(\begin{bmatrix}JW_{0}\\ I_{k}\end{bmatrix}-\begin{bmatrix}JW_{0}\\ 0\end{bmatrix}\biggr)=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}$.
Repeating this procedure with $\dot{Y}_{\rho}\mathrm{e}^{tJW_{0}}$, we obtain
$\ddot{Y}_{\rho}\mathrm{e}^{tJW_{0}}+\dot{Y}_{\rho}\mathrm{e}^{tJW_{0}}JW_{0}=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}$.
@@
$\ddot{Y}_{\rho}\mathrm{e}^{tJW_{0}}=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}-JV_{0}\\ 0\end{bmatrix}$. (36)
@@
$\dot{Y}_{\rho}^{T}A\dot{Y}_{\rho}=$ $\left(\mathrm{e}^{-tJW_{0}}\right)^{T}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}J&W_{0}\\ -W_{0}&V_{0}\end{bmatrix}\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}[0\;\;I_{k}]\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=$ $\,\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}},$ (37)
@@
Finally, left multiplying (37) by $Y_{\rho}J$ yields
$Y_{\rho}J(\dot{Y}_{\rho}^{T}A\dot{Y}_{\rho})=$ $\,Y_{\rho}J\mathrm{e}^{-tW_{0}^{T}J}V_{0}\mathrm{e}^{-tJW_{0}}$
$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}}J^{2}\mathrm{e}^{tJW_{0}}JV_{0}\mathrm{e}^{-tJW_{0}}$
$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}JV_{0}\mathrm{e}^{-tJW_{0}}$
$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}JV_{0}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}},$ (38)
@@
$(\mathrm{e}^{-tJW_{0}})^{T}=\mathrm{e}^{-tW_{0}^{T}J}=J^{2}\mathrm{e}^{-tW_{0}^{T}J}=J\mathrm{e}^{tJW_{0}}J.$
@@
Finally, the uniqueness of the solution is derived from that of the problems (30) and (34) which is a well-known fact in the theory of linear differential equations in Banach spaces, see, e.g., (Deimling1977, , Sect. 1.1). ∎
The second part of the proof for Proposition 4, implicitly assumes that the curve $Y_{\rho}(t)$ in (33) is feasible. The statement below confirms it.
@@
$Y_{\rho}^{T}AY_{\rho}=\left(\mathrm{e}^{-tJW_{0}}\right)^{T}[I_{k}\;\;0]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}}.$ (39)
@@
$\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)=\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi=\begin{bmatrix}J&W_{0}\\ -W_{0}&V_{0}\end{bmatrix}$.
@@
$Y_{\rho}^{T}AY_{\rho}=$ $\left(\mathrm{e}^{-tJW_{0}}\right)^{T}[I_{k}\;\;0]\begin{bmatrix}J&W_{0}\\ -W_{0}&V_{0}\end{bmatrix}\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJW_{0}}$
$=$ $J\mathrm{e}^{tJW_{0}}J^{2}\mathrm{e}^{-tJW_{0}}=J,$
in which the second identity follows from the fact that $\left(\mathrm{e}^{-tJW_{0}}\right)^{T}=J\mathrm{e}^{tJW_{0}}J.$
@@
Given $X\in\mathrm{St}_{A}(k,n)$, $Z\in T_{X}\mathrm{St}_{A}(k,n)$ and set $W_{0}=X^{T}AZ$, $V_{0}=Z^{T}AZ$. The curve
$Y(t)=[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}W_{0}&-V_{0}\\ I_{k}&W_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tW_{0}}$ (40)
@@
Restricting to the orthogonal Stiefel manifold, the quasi-geodesic (40) reduces to the geodesic obtained in (EdelAS98, , Subsect. 2.2). It is worth to note that the one in (EdelAS98, , Subsect. 2.2) is indeed associated with the Euclidean metric. Nevertheless, by direct verification, one can show that the normal space, orthogonal projections onto the tangent and normal spaces under the Euclidean and under the canonical metrics are identical. Therefore, for the orthogonal Stiefel manifold, the embedded geodesic associated with the Euclidean metric is also the quasi-geodesic associated with the canonical metric. Hence, (33) is a generalization of this quasi-geodesic to the indefinite Stiefel manifold.

### v2

@@ 4 Quasi-geodesics with respect to the generalized canonical metric
A geodesic on a manifold is a smooth curve with vanishing acceleration. In the dawn of Riemannian optimization, the solution is updated by searching along the geodesic which goes through the current solution in an appropriate tangent direction EdelAS98 ; Gaba82 ; Smit94 . In the Euclidean space, zero acceleration means $\ddot{X}(t)=0$, or equivalently $X(t)=\dot{X}(0)t+X(0)$, which yields a straight line. That is, geodesics can be viewed as a generalization of lines in the Euclidean space to the Riemannian manifold. In a general setting, the notion of acceleration depends on the Riemannian connection (AbsiMS08, , Sect. 5.4) and therefore the differential equation determining geodesics is usually much more involved. This makes the task of finding an explicit formula for the geodesic, given initial information $X(0),\dot{X}(0)$, very difficult. Fortunately, in optimization, only local information encoded by geodesics and its feasibility are interesting due to their usage in updating the iterate. Thus, one usually retains these features in approximating geodesics which results in the so-called quasi-geodesics; see, e.g., JurdML2020 ; KrakMLB2017 ; NishA2005 for more approaches to the quasi-geodesics on the orthogonal Stiefel manifold.
In this paper, we follow a procedure similar to that of (GSAS21, , Subsect. 4.3), which was adopted from (EdelAS98, , Subsect. 2.2). Namely, we replace the acceleration that should be formulated in accordance with the Riemannian connection associated with the proposed generalized canonical metric by the one corresponding to the Euclidean metric, which turns out to be $\ddot{Y}(t).$ In view of the normal vectors formulated in Proposition 1, we define a quasi-geodesic with respect to the generalized canonical metric $g_{\rho,X_{\perp}}$ to be any solution $Y_{\rho}(t)\in\mathrm{iSt}(k,n)$ of the initial value problem
@@
with some $U(t)$ such that $JU(t)\in\mathrm{Sym}(k)$ for all $t\geq 0$, although the obtained result apparently holds for $t<0$ too. In this section, a closed-form solution to this equation is presented. For the sake of brevity, we refrain from writing the variable $t$ in our presentation except for the arguments in which the dependence on $t$ is highlighted.
@@
$(-YU)^{T}AY+2\dot{Y}^{T}A\dot{Y}+Y^{T}A(-YU)=0$
or $\dot{Y}^{T}A\dot{Y}=JU$ due to the symmetry of $JU$. Hence, $U=J\dot{Y}^{T}A\dot{Y}$ and therefore, the equation (25) becomes
@@
Next, let us set $S\coloneqq Y^{T}A\dot{Y}$ and $V\coloneqq\dot{Y}^{T}A\dot{Y}$. Thanks to the relation (26), $S\in\mathrm{Skew}(k)$ for $t\geq 0.$ Moreover, using (28), we get
$\dot{S}=V+Y^{T}A\ddot{Y}=V-Y^{T}AYJV=V-V=0,\mbox{ for }t\geq 0.$
Hence, $S(t)=S(0)\eqqcolon S_{0}\in\mathrm{Skew}(k)$ for $t\geq 0$. We proceed to determine $V$ which is symmetric. Taking the derivative of $V$ with (28) in mind, we have that
@@
$\dot{V}=S_{0}JV-VJS_{0},\ V(0)=V_{0},$ (29)
admits a solution $V(t)=\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}}.$ Indeed, differentiating such $V(t)$ yields
$\dot{V}$ $=-S_{0}^{T}J\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}}-\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}}JS_{0}$
$=-S_{0}^{T}JV-VJS_{0}=S_{0}JV-VJS_{0}.$
Next, we present a technical lemma, which is similar to (GSAS21, , Prop. 4.6).
@@
$Y_{\rho}(t)=Y_{\rho}^{\mathrm{qgeo}}(t,Z)\coloneqq[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}JS_{0}&-JV_{0}\\ I_{k}&JS_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}},$ (32)
where $S_{0}=Y_{\rho}^{T}(0)A\dot{Y}_{\rho}(0)=X^{T}AZ$ and $V_{0}\coloneq\dot{Y}_{\rho}^{T}(0)A\dot{Y}_{\rho}(0)=Z^{T}AZ.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(Y_{\rho}(t)\mathrm{e}^{tJS_{0}}\right)$ $=\dot{Y_{\rho}}(t)\mathrm{e}^{tJS_{0}}+Y_{\rho}(t)\mathrm{e}^{tJS_{0}}JS_{0}.$
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}\right)$ $=\ddot{Y_{\rho}}(t)\mathrm{e}^{tJS_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}JS_{0}$
$=\left(-Y_{\rho}(t)JV\right)\mathrm{e}^{tJS_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}JS_{0}$
$=-Y_{\rho}(t)J\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}}\mathrm{e}^{tJS_{0}}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}JS_{0}$
$=-Y_{\rho}(t)\mathrm{e}^{tJS_{0}}JV_{0}+\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}JS_{0},$
in which we have invoked (29) to replace $V\coloneq\dot{Y_{\rho}}^{T}A\dot{Y_{\rho}}$ by its explicit form $V=\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}}$ when $Y_{\rho}(t)\in\mathrm{iSt}(k,n),$ for all $t\geq 0.$ Therefore,
$\frac{\mathrm{d}}{\mathrm{d}t}\left[Y_{\rho}(t)\mathrm{e}^{tJS_{0}}\;\;\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}\right]=\left[Y_{\rho}(t)\mathrm{e}^{tJS_{0}}\;\;\dot{Y}_{\rho}(t)\mathrm{e}^{tJS_{0}}\right]\begin{bmatrix}JS_{0}&-JV_{0}\\ I_{k}&JS_{0}\end{bmatrix}$. (33)
@@
$Y_{\rho}(t)$ $=\left[Y_{\rho}(0)\;\;\dot{Y}_{\rho}(0)\right]\mathrm{exp}\left(t\begin{bmatrix}JS_{0}&-JV_{0}\\ I_{k}&JS_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
$=[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}JS_{0}&-JV_{0}\\ I_{k}&JS_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}},$
@@
Next, we will show that this is indeed a global solution in $\mathrm{iSt}(k,n)$ to the problem (31). For future reference, we present the feasibility part in Proposition 5 below. In the interest of brevity, we set $\Psi\coloneq\left[\begin{smallmatrix}JS_{0}&-JV_{0}\\ I_{k}&JS_{0}\end{smallmatrix}\right]$. In view of (32), it holds that
$\dot{Y}_{\rho}\mathrm{e}^{tJS_{0}}+Y_{\rho}\mathrm{e}^{tJS_{0}}JS_{0}=\frac{\mathrm{d}}{\mathrm{d}t}\left(Y_{\rho}\mathrm{e}^{tJS_{0}}\right)=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}I_{k}\\ 0\end{bmatrix}$.
@@
$\dot{Y}_{\rho}\mathrm{e}^{tJS_{0}}=$ $[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}I_{k}\\ 0\end{bmatrix}-[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}JS_{0}$ (34)
$=$ $[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\biggr(\begin{bmatrix}JS_{0}\\ I_{k}\end{bmatrix}-\begin{bmatrix}JS_{0}\\ 0\end{bmatrix}\biggr)=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}$.
Repeating this procedure with $\dot{Y}_{\rho}\mathrm{e}^{tJS_{0}}$, we obtain
$\ddot{Y}_{\rho}\mathrm{e}^{tJS_{0}}+\dot{Y}_{\rho}\mathrm{e}^{tJS_{0}}JS_{0}=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}$.
@@
$\ddot{Y}_{\rho}\mathrm{e}^{tJS_{0}}=[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}-JV_{0}\\ 0\end{bmatrix}$. (35)
@@
$\dot{Y}_{\rho}^{T}A\dot{Y}_{\rho}=$ $\left(\mathrm{e}^{-tJS_{0}}\right)^{T}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
$=$ $\,\mathrm{e}^{-tS_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}J&S_{0}\\ -S_{0}&V_{0}\end{bmatrix}\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
$=$ $\,\mathrm{e}^{-tS_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
$=$ $\,\mathrm{e}^{-tS_{0}^{T}J}[0\;\;I_{k}]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\mathrm{exp}\left(t\Psi\right)\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
$=$ $\,\mathrm{e}^{-tS_{0}^{T}J}[0\;\;I_{k}]\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi\begin{bmatrix}0\\ I_{k}\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
$=$ $\,\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}},$ (36)
@@
At the end, left multiplying (36) by $Y_{\rho}J$ yields
$Y_{\rho}J(\dot{Y}_{\rho}^{T}A\dot{Y}_{\rho})=$ $\,Y_{\rho}J\mathrm{e}^{-tS_{0}^{T}J}V_{0}\mathrm{e}^{-tJS_{0}}$
$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}}J^{2}\mathrm{e}^{tJS_{0}}JV_{0}\mathrm{e}^{-tJS_{0}}$
$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}JV_{0}\mathrm{e}^{-tJS_{0}}$
$=$ $\,[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}JV_{0}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}},$ (37)
@@
$(\mathrm{e}^{-tJS_{0}})^{T}=\mathrm{e}^{-tS_{0}^{T}J}=J^{2}\mathrm{e}^{-tS_{0}^{T}J}=J\mathrm{e}^{tJS_{0}}J.$ (38)
@@
The uniqueness of the solution is derived from that of the problems (29) and (33) which is a well-known fact in the theory of linear differential equations in Banach spaces; see, e.g., (Deimling1977, , Sect. 1.1).
@@
$Y_{\rho}^{T}AY_{\rho}=\left(\mathrm{e}^{-tJS_{0}}\right)^{T}[I_{k}\;\;0]\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}}.$ (39)
@@
$\left(\mathrm{exp}(t\Psi)\right)^{T}\begin{bmatrix}X^{T}\\ Z^{T}\end{bmatrix}A[X\;\;Z]\mathrm{exp}\left(t\Psi\right)=\begin{bmatrix}0&J\\ -J&0\end{bmatrix}\Psi=\begin{bmatrix}J&S_{0}\\ -S_{0}&V_{0}\end{bmatrix}$.
@@
$Y_{\rho}^{T}AY_{\rho}=$ $\left(\mathrm{e}^{-tJS_{0}}\right)^{T}[I_{k}\;\;0]\begin{bmatrix}J&S_{0}\\ -S_{0}&V_{0}\end{bmatrix}\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tJS_{0}}$
$=$ $J\mathrm{e}^{tJS_{0}}J^{2}\mathrm{e}^{-tJS_{0}}=J,$
in which the second identity follows from the fact that $\left(\mathrm{e}^{-tJS_{0}}\right)^{T}=J\mathrm{e}^{tJS_{0}}J$ derived in (38). ∎
@@
Given $X\in\mathrm{St}_{A}(k,n)$, $Z\in T_{X}\mathrm{St}_{A}(k,n)$, and set $S_{0}=X^{T}AZ$, $V_{0}=Z^{T}AZ$. The curve
$Y(t)=[X\;\;Z]\mathrm{exp}\left(t\begin{bmatrix}S_{0}&-V_{0}\\ I_{k}&S_{0}\end{bmatrix}\right)\begin{bmatrix}I_{k}\\ 0\end{bmatrix}\mathrm{e}^{-tS_{0}}$ (40)
@@
Restricting to the orthogonal Stiefel manifold, the quasi-geodesic (40) reduces to the geodesic obtained in (EdelAS98, , Subsect. 2.2). It is worth noting that the one in (EdelAS98, , Subsect. 2.2) is indeed associated with the Euclidean metric. Nevertheless, by direct verification, one can show that the normal space, orthogonal projections onto the tangent and normal spaces under the Euclidean and under the canonical metrics are identical. Therefore, for the orthogonal Stiefel manifold, the embedded geodesic associated with the Euclidean metric is also the quasi-geodesic associated with the canonical metric. Hence, (32) is a generalization of this quasi-geodesic to the indefinite Stiefel manifold.